OpenStudy (anonymous):

An auto technician has 12 ounces of battery acid that is 75% acid. How many ounces of water must the technician add to make a solution that is 25% acid?

OpenStudy (gorv):

75 % of 12 = 9 ounce

OpenStudy (gorv):

and 3 ounce of water

OpenStudy (gorv):

new total = 12+x water = 3+x acid = 9

OpenStudy (gorv):

\[\frac{ 9 }{ 12+x }=\frac{ 25 }{ 100 }\]

OpenStudy (gorv):

\[\frac{ 9 }{ 12+x }=\frac{ 1 }{ 4 }\]

OpenStudy (gorv):

9*4=12+x 36=12+x x=24